OpenStudy (anonymous):

i need help in solving this 0.8+0.2(y-1)=1.8

OpenStudy (anonymous):

I think step 1 should be to expand / distribute the bracket: a(b+c) = a*b + a*c \(\Rightarrow\) 0.2(y-1) = ?

OpenStudy (anonymous):

if you get stuck multiply everything by 5

OpenStudy (anonymous):

maybe you will see it better

OpenStudy (anonymous):

4+y-1=9 y=6